OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −4(3d − 1) + 4d

OpenStudy (unklerhaukus):

\[−4(3d − 1) + 4d \] \[=(−4\times3d) +(-4\times− 1) + 4d \] \[=\]